# Anomalous Criticality of Absorbing State Transition toward Jamming

### v1

@@ V Conclusion
In summary, we investigate the critical behavior of the BRO model in dense monodisperse and binary systems in $d=2$ to 4. We find that anomalous critical phenomena emerge as the critical density increases by decreasing $\varepsilon$. For 3D monodispersed systems at intermediate critical density, the crystallization interrupts the dynamic phase transition. For binary systems at intermediate critical density, the measured critical exponents deviate from the Manna universality class due to the caging effects, signifying the existence of new universality of dynamic phase transition, i.e., an absorbing to active glass transition. For both monodispersed and binary systems at higher density, the quenched disorder of contact networks triggers pronounced Griffiths effect, smearing out the dynamic phase transition. In the $\varepsilon\rightarrow 0$ limit, the BRO model becomes equivalent to DP on heterogeneous network. Our results demonstrate that although the BRO model can generate RCP configurations, the criticality of the dynamic transition is fundamentally reshaped by the geometric rigidity of the RCP state. These findings discourage previous attempts to directly associate the jamming transition with the Manna universality class. We also propose a field theory with fractional time dynamics which unifies both the absorbing to actve-glass transition and heterogeneous DP, revealing a deep connection between dynamic criticality and spatial annealed/quenched disorder. Our work could potentially be related to ultra-stable glass and the Gardner transition [61, 62, 63, 64] and open a door to study the dynamic criticality of more complicated artificial neuron networks [36], like the learning process of spherical negative perceptron model [31].
@@ VI Method
where $n_{a}(i)$ represents the number of infected neighbors of the particle $i$, and $n_{neib}(i)$ is the total number of neighbors of the particle $i$. The recovery rate of each site, i.e., the probability that an infected site becomes healthy in the next time step, is $P_{r}=1-P_{e}$. We establish a contact network based on the particle configuration generated by BRO dynamics, where the positions of particles are the nodes of the network. A contact bond is assumed between the two nodes if their distance is less than $0.5\sigma$. We introduce the heterogeneity by making the infection rate of site $i$ undermined by the activity duration $f_{ad}(i)$ of the particle, i.e.,
@@
Here $f_{ad}(i)$ is in the range of $(0,1)$. As $k$ increases, the heterogeneity becomes more prominent. We calculated the data of changing the basic infection rate $P_{e}$ when k=1, 2 and 4. We find that the Griffith effect become stronger as $k$ increases (see Fig. S11).
@@ Acknowledgments:
The authors are grateful to Ran Ni, Ning Xu, Hao Hu and Hua Tong for helpful discussion. This work is supported by the National Natural Science Foundation of China (No. 12347102, No. 12275127), the Innovation Program for Quantum Science and Technology (No. 2024ZD0300101), the Natural Science Foundation of Jiangsu Province (No. BK20250058, No. BK20233001), Program for Innovative Talents and Entrepreneur in Jiangsu, the Fundamental Research Funds for the Central Universities (0204-14380249, KG202501). The simulations are performed on the High-Performance Computing Center of Collaborative Innovation Center of Advanced Microstructures, the High-Performance Computing Center (HPCC) of Nanjing University.

### v2

@@ V Conclusion
In summary, we investigate the critical behavior of the BRO model in dense monodisperse and binary systems in $d=2$ to 4. We find that anomalous critical phenomena emerge as the critical density increases by decreasing $\varepsilon$. For 3D monodispersed systems at intermediate critical density, the crystallization interrupts the dynamic phase transition. For binary systems at intermediate critical density, the measured critical exponents deviate from the Manna universality class due to the caging effects, signifying the existence of new universality of dynamic phase transition, i.e., an absorbing to active glass transition. For both monodispersed and binary systems at higher density, the quenched disorder of contact networks triggers the pronounced Griffiths effect, smearing out the dynamic phase transition. In the $\varepsilon\rightarrow 0$ limit, the BRO model becomes equivalent to DP on heterogeneous network. Our results demonstrate that although the BRO model can generate RCP configurations, the criticality of the dynamic transition is fundamentally reshaped by the geometric rigidity of the RCP state. These findings discourage previous attempts to directly associate the jamming transition with the Manna universality class. We also propose a field theory with fractional time dynamics which unifies both the absorbing to active-glass transition and heterogeneous DP, revealing a deep connection between dynamic criticality and spatial annealed/quenched disorder. Our work could potentially be related to ultra-stable glass and the Gardner transition [64, 65, 66, 67, 36] and opens the door to studying the dynamic criticality of more complicated artificial neural networks [39], like the learning process of spherical negative perceptron model [32].
@@ VI Method
where $n_{a}(i)$ represents the number of infected neighbors of the particle $i$, and $n_{neib}(i)$ is the total number of neighbors of the particle $i$. The recovery rate of each site, i.e., the probability that an infected site becomes healthy in the next time step, is $P_{r}=1-P_{e}$. We establish a contact network based on the particle configuration generated by BRO dynamics, where the positions of particles are the nodes of the network. A contact bond is assumed between the two nodes if their distance is less than $0.5\sigma$. We introduce the heterogeneity by making the infection rate of site $i$ coupling with the activity duration $f_{ad}(i)$ of the particle, i.e.,
@@
Here $f_{ad}(i)$ is in the range of $(0,1)$. As $k$ increases, the heterogeneity becomes more prominent. We study different basic infection rate cases under kk=1, 2 and 4. We find that the Griffith effect becomes stronger as $k$ increases (see Fig. S10).
@@ Acknowledgments:
The authors are grateful to Ran Ni, Ning Xu, Hao Hu and Hua Tong for helpful discussions. This work is supported by the National Natural Science Foundation of China (No. 12347102, No. 12275127), the Innovation Program for Quantum Science and Technology (No. 2024ZD0300101), the Natural Science Foundation of Jiangsu Province (No. BK20250058, No. BK20233001), Program for Innovative Talents and Entrepreneurs in Jiangsu, the Fundamental Research Funds for the Central Universities (0204-14380249, KG202501). The simulations are performed on the High-Performance Computing Center of Collaborative Innovation Center of Advanced Microstructures, the High-Performance Computing Center (HPCC) of Nanjing University.